Last month Molly earned $200 a week and got a bonus of $150. During the same period, Harry earned $300 a week and got a bonus of $50. If these earnings continue, when will their earnings be the same, and what will the amount be?
Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

i'll keep this simple and assume that there are 4 weeks in a month.
the actual figure is more like 13/3 weeks in a month.
this comes out to approximately 4.33 weeks in a month, rounded to 2 decimal places.
asume 4 weeks in a month for simplicity.

if you're talking about their earnings being the same in the same number of months, that won't happen.
if you're talking about their earnings being the same after molly works x months and harry works y months, then you can figure it out.

you can't however, solve for a specific month.
the answer will be relative to one of the variable.
i'll go through it and you'll see what i mean.

for molly, she earns 200 per week and gets a bonus of 150 at the end of the month.
her monthly earnings will therefore be 4 * 200 = 800 + 150 = 950
for harry, he earns 300 per week and gets a bonus of 50 at the end of the month.
his monthly earnings will therefore be 4 * 300 = 1200 + 50 = 1250.

you can see that their monthly earnings will never be the same in the same number of months for each.
harry will always make more than her regardless of the number of months.

let x equal the number of months molly works and y equal the number of months harry works.
their earnings will be the same when 950 * x = 1250 * y.
solve for y to get y = 950 * x / 1250
this comes out to be y = .76 * x

what this says is that, if molly works x months, then their total earnings will be the same when harry works .76 * x months.

if molly works x = 12 months, then their total earnings will be the same when harry works .76 * 12 = 9.12 months.

molly's total earnings will be 12 * 950 = 11,400
harry's total earnings will be .76 * 12 * 1250 = 11,400

this formula will work no matter how many months molly works.

for example, if molly works 75 months, then their earnings will be the same when harry works .76 * 75 = 57 months.

molly's total earnings will be 75 * 950 = 71250
harry's total earnings will be 57 * 1250 = 71250



Answer by MathTherapy(10552)


Last month Molly earned $200 a week and got a bonus of $150. During the same period, Harry earned $300 a week and got a bonus of $50. If these earnings continue, when will their earnings be the same, and what will the amount be?
Am I missing something here? Weren't their earnings the same during a week last month? 
Didn't Molly earn $350 (200 + 150) during a week last month?
Didn't Harry earn $350 (300 + 50) during a week last month?
Now, aren't those amounts the same?